the mean score of random sample of 22 student applicants who took the college entrance examination is calculated to be 76.9 the population standard deviation is known to be 27.4 find the 99% confidence interval for the mean of the entire student applicants (hint: ESTIMATION OF PARAMETERS)


1
Expert's answer
2022-04-07T03:33:57-0400

"99\%CI=(\bar x-z_{0.005}\frac{\sigma}{\sqrt{n}},\bar x+z_{0.005}\frac{\sigma}{\sqrt{n}})="

"=(76.9-2.33\frac{27.4}{\sqrt{22}},76.9+2.33\frac{27.4}{\sqrt{22}})=(63.29,90.51)."